Bovine Company, a wholesale distributor of umbrellas, has been experiencing losses for some time, as shown by its most recent monthly contribution format income statement: $2,190,000 Sales Variable 887,520 expenses Contribution Fixed expenses 1,343,000 Operating loss $ (40,520) 1,302,480 margin In an effort to isolate the problem, the president has asked for an income statement segmented by geographic market. Accordingly, the Accounting Department has developed the following Geographic Market North Central 619,000 $818,000 $753,000 South Sales Variable expenses as a 50% 32% 42% percentage of sales Traceable fixed expenses $350,000 $528,000 319,000 Required: 1. Prepare a contribution format income statement segmented by geographic market, as requested by the president.
Answer #1

Answer:-1)-

Bovine Company
Income Statement (Contribution Format)
Divisons
Total Company North South South
Particulars Amonut Amonut Amonut Amonut
$ $ $ $
Sales 2190000 619000 818000 753000
Less:- Variable expenses 887520 619000*50%=309500 818000*32%=261760 753000*42%=316260
Contribution Margin 1302480 309500 556240 436740
Less:-Traceble fixed expenses 1197000 350000 528000 319000
Territorial segment margin 105480 -40500 28240 117740
Less:-Common fixed expenses 1343000-1197000=146000
Net opreating income -40520

2a)-

Incremental Operating Income
Particulars Amount
$
Incremental sales $818000*10% 81800
Contrbution margin ratio 68% $81800*68% 55624
Incremental contribution margin 55624
Less:- Incremental advertising expenses 44000
Incremental operating income 11624

2b)-Yes, the advertising program should be initiated.